OpenStudy (anonymous):

How many numbers of 4 digits each greater than 5000 can be formed from the digits 1,2,3,4,5,6,7?..help plzz..

OpenStudy (agent0smith):

Three choices for the first digit (5, 6 or 7 are the only ones to make 4 digit numbers > 5000) then 7 choices for the remaining digits, so it should be... 3*7*7*7 = ?

OpenStudy (anonymous):

1029??

OpenStudy (agent0smith):

@Luis_Rivera what...? Not sure what you meant, but 0 isn't an choice, so all 5xxx are greater than 5000.

OpenStudy (agent0smith):

Smallest number we can make with 1,2,3,4,5,6,7 which is greater than 5000 is 5111, so there's 3 options, 5 6 and 7 for the first digit.

OpenStudy (agent0smith):

uh, why not...?

OpenStudy (agent0smith):

@Luis_Rivera "How many numbers of 4 digits each greater than 5000 can be formed from the digits 1,2,3,4,5,6,7?" Here, 5xxx is greater than 5000... i'm not following you...

OpenStudy (anonymous):

so what are the answer?

OpenStudy (anonymous):

is*

OpenStudy (agent0smith):

\[ \large 3*7^3\]

OpenStudy (anonymous):

if without repeating numbers??

OpenStudy (agent0smith):

W/o repeating it'll be... 3*6*5*4

OpenStudy (kropot72):

Taking 5 as the first digit there are 6P3 permutations of the remaining 6 digits. Taking 6 as the first digit there are 6P3 permutations of the remaining 6 digits. Taking 7 as the first digit there are 6P3 permutations of the remaining 6 digits. Total number of permutations are \[3\times 6P3=3\times \frac{6!}{3!}=360\]

OpenStudy (agent0smith):

^correct, for no repetition.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

without repeating=360..repeating - 28800???....3x6P3x5p2x4p1?

OpenStudy (agent0smith):

^ no, with repeating it's 3*7*7*7 = 3*7^3 = 1029